OpenStudy (anonymous):

please heeellppp... a box holds 20 cards numbered 1-20. you select a card. without replacing the first card you select a second. find the probability of getting a 1, then 20. steps please.

OpenStudy (anonymous):

\[\frac{1}{20}\times \frac{1}{19}\]

OpenStudy (anonymous):

oh wow thanks i got i right